OpenStudy (anonymous):

What is the value of x in the equation −x = 4 − 3x + 6? 5 10 −5 −10

OpenStudy (batman19991):

1.5(x + 4) – 3 = 4.5(x – 2) 1.5x + 6 - 3 = 4.5x - 9 3 + 9 = 4.5x - 1.5x 12 = 3x x = 12/3 x = 4 so the closest answer to this would be 5 i think i dont really know

OpenStudy (anonymous):

Which statement is true about the equation fraction 3 over 4z = fraction 1 over 4z − 3 + 5? It has no solution. It has one solution It has two solutions. It has infinitely many solutions.

OpenStudy (batman19991):

i dont know i dont think it has a solution

OpenStudy (anonymous):

K THXS!!!!!!

OpenStudy (batman19991):

no problem let me know if im wrong